A condenser allow an enthalpy of 2280 kJ/kg into it and exits at 163 kJ/kg at a rate of4500kg/hour. The condenser cooling water increase at a rate of 148000kJ/mm. what is the rate at which the condenser flows to the atmosphere
1
Expert's answer
2021-08-10T18:00:14-0400

We know that

"\frac{dQ}{dt}=m_s(h_1-h_2)"

"\frac{dQ}{dt}=\frac{4500}{3600}\times(2280-163)=2.64kj/sec"